OpenStudy (mralexrest):

Factor Completely,9n^2+30n+25

OpenStudy (anonymous):

since the original function can be expressed as 9n^2+2*15n+25.

OpenStudy (anonymous):

and you can see that 9= 3*3

OpenStudy (anonymous):

25=5*5. it is pretty obvious that 3*5=15 which is the middle term you need.